OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity 4 times x all over y all to the second power. I feeel like I will get this wrong.

OpenStudy (mertsj):

\[(\frac{4x}{y})^2=\frac{4^2x^2}{y^2}=\frac{16x^2}{y^2}\]